OpenStudy (anonymous):

solve for y: -6y - 1 +7(y +1) = 4y -2

OpenStudy (anonymous):

y= 8/3

OpenStudy (anonymous):

thanks, but I need to have work shown

OpenStudy (anonymous):

-6y - 1 +7(y +1) = 4y -2 -6y - 1 +7y + 7 = 4y -2 (Combine like terms) y + 6 = 4y - 2 6= 3y + 2 8= 3y 8/3 = y y= 8/3 <--- Final Answer

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Welcome